40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 12 20:15:22 2016, Lydia, Update, SUS, Diagonalization in air ITMY_fit.jpgBS_diag.jpg
    Reply  Tue Sep 13 19:18:43 2016, Lydia, Update, SUS, Diagonalization in air ITMY_diag.jpg38.png
       Reply  Wed Sep 14 20:27:03 2016, Lydia, Update, SUS, Diagonalization ETMY_diag.png
          Reply  Thu Sep 15 18:37:20 2016, Lydia, Update, SUS, Diagonalization ETMY_osemspec.pngITMY_osemspec.png
             Reply  Fri Sep 16 19:14:27 2016, Lydia, Update, SUS, Diagonalization ETMY_yawdrivedefault.pngETMY_yawdrivenew.pngETMY_yawdriveold.png57.png
Message ID: 12484     Entry time: Mon Sep 12 20:15:22 2016     Reply to this: 12490
Author: Lydia 
Type: Update 
Category: SUS 
Subject: Diagonalization in air 

[Lydia, Teng]

We ran the scripts to diagonalize the damping matrices using the free swinging data from staurday night/sunday morning. The actual entries used for damping have not been changed. However, we did generate updated matrices for all the main optics (not including the mode cleaner optics, which were not free swinging over the weekend).

  • The scripts appear to be mostly working as intended, with a couple of issues:
    • The plots made by makeSUSSpectra claim to be showing spectra of the individual OSEM readings, but are actually dofs calculated using the ideal input matrix.
    • The existing parameters file (for the peak finding) was only fitting the lorentz peaks to a very narrow band of data, close to the bandwidth of the spectrum. Too narrow a band means that the initial guess must be very close, and also means there are not enough points to fit to.
      • We modified a copy of the paramters file to use a wider band (~.1 Hz) for fitting, and also use updated estimates of the mode frequencies.
      • This was largely successful, but the ITMY POS peak is very close to the SIDE peak, and POS is also stringly coupled to SIDE, so the wider bandwidth fitting can't separate the peaks. (See attachment 1)
      • A longer time series, plus more accurate initial guesses for the resonance frequencies, would allow us to fit to a smaller (~.03 Hz) band without encountering the stated issues.
      • A better way than manually examining plots to choose an initial frequency guess would be to automatically start at the overall maximum point in the spectrum between 0.4 and 1.5 Hz
  • Most of the diagonalization results seem good: "Badness" numbers of 4-6 and secondary peaks very supressed or absent on spectra plotted in dof basis (See attachment 2). ITMY, perhaps beacuse of a related issue, has phase problems with the matrix elements that result in messages like "osem/dof 2/1 is imaginary."
Attachment 1: ITMY_fit.jpg  101 kB  | Hide | Hide all
ITMY_fit.jpg
Attachment 2: BS_diag.jpg  202 kB  | Hide | Hide all
BS_diag.jpg
ELOG V3.1.3-